INFRARED SPECTRAL STUDIES OF THE THERMALLY-DRIVEN CHEMISTRY PRESENT ON ICY SATELLITES

\underline{MARK J. LOEFFLER}, REGGIE L. HUDSON, NASA Goddard Space Flight Center, Astrochemistry Laboratory, Mail Code 691.1, Greenbelt, Md 20771.

Remote sensing of Jupiters icy satellites has revealed that even though their surfaces are composed mostly of water ice, molecules such as SO_{2}, CO_{2}, H_{2}O_{2}, O_{2}, and O_{3} also are present. On Europa, a high radiation flux is believed to play a role in the formation of many of the minor species detected, and numerous laboratory studies have been devoted to explore this hypothesis. In this presentation we will discuss some of our recent research on another alteration pathway, thermallydriven chemical reactions, which are also important for understanding the chemical evolution of Europa’s surface and sub-surface ices. We will focus on the infrared spectra of and reactions between H_{2}O, SO_{2}, and H_{2}O_{2} at 80 - 130 K.